OpenStudy (anonymous):

Solve the Equation 1/(x-1)+6/(x+1)=8/(x^2-1)

OpenStudy (luigi0210):

Get common denominators first

OpenStudy (anonymous):

how do i find the common denominator for this type of problem?

OpenStudy (luigi0210):

Factor out the denominators and see what they have on common and what they are missing

OpenStudy (anonymous):

so (x-1}{x+1)(x^2-1) is the common denominator? right

OpenStudy (luigi0210):

Nope, you can factor that (x^2-1) you know

OpenStudy (jdoe0001):

keep in mind that \(\bf x^2-1 \implies x^2-1^2\)

OpenStudy (jdoe0001):

and that \(\bf (a-b)(a+b) = (a^2-b^2)\)

OpenStudy (luigi0210):

@jdoe0001 You wanna handle this?

OpenStudy (jdoe0001):

ok

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{x-1}+\cfrac{6}{x+1}=\cfrac{8}{x^2-1} \implies \cfrac{1}{x-1}+\cfrac{6}{x+1}=\cfrac{8}{x^2-1^2}\\ \implies \cfrac{1}{x-1}+\cfrac{6}{x+1}=\cfrac{8}{\square?}\)

OpenStudy (jdoe0001):

so what do you think about \(\bf x^2-1^2\) ?

OpenStudy (anonymous):

Got it, make sense, thanks. here is another one if you can help me with this. 16x-7/(8x+3)=2-3/x